OpenStudy (asdasdasd):

Reallly need Help!!! Will give medal. How can long division and synthetic division be used to find zeros of polynomial functions?

OpenStudy (asdasdasd):

@lucaz Can you please help me?

OpenStudy (lucaz):

long division helps you to factor the polynomial, you get something like f(x)=(x+a)(x-b)(x+c) the roots would be -a, b and -c

OpenStudy (asdasdasd):

Okay

OpenStudy (lucaz):

for example f(x)=x^3+5x^2-2x-24 factoring by long division you get f(x)=(x+3)(x-2)(x+4)

OpenStudy (asdasdasd):

But isn't long division like x2–2x+4 by x–6.

OpenStudy (asdasdasd):

okay

OpenStudy (lucaz):

yes that's long division..

OpenStudy (asdasdasd):

x^2-2x+4*

OpenStudy (lucaz):

take a look here http://www.purplemath.com/modules/solvpoly.htm

OpenStudy (asdasdasd):

Okay i understand now.

OpenStudy (asdasdasd):

How about synthetic division?

OpenStudy (lucaz):

they have the same goal, the operation is just a bit different

OpenStudy (lucaz):

you find the same result

OpenStudy (asdasdasd):

Can you give me an example of synthetic?